# Learning out-of-time-ordered correlators with classical kernel methods

### v1

@@ C.6 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{3}$
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 5.2940 5.8909 6.0765 6.1494 6.1253 6.0140 5.8459 5.6489 MAE 4.2697 4.8224 4.9492 4.9773 4.9052 4.7378 4.5049 4.2420 $R^{2}$1.8253 -2.0274 -1.9619 -1.8855 -1.7631 -1.5998 -1.4225 -1.2536 Polynomial RMSE 1.1585 0.9611 0.9857 0.9956 0.9778 1.0404 1.0888 1.1037 MAE 0.8594 0.7375 0.7432 0.7518 0.7466 0.7980 0.8239 0.8420 $R^{2}$ 0.8647 0.9194 0.9221 0.9244 0.9296 0.9222 0.9160 0.9140 RBF RMSE 0.3834 0.3908 0.3831 0.6243 0.5514 0.5142 0.4808 0.4614 MAE 0.2247 0.2704 0.2342 0.4376 0.3829 0.3485 0.3089 0.2831 $R^{2}$ 0.9852 0.9867 0.9882 0.9703 0.9776 0.9810 0.9836 0.9850 Laplacian RMSE 0.6000 0.6719 0.6060 0.5842 0.5770 0.5740 0.5768 0.5884 MAE 0.3753 0.4879 0.4285 0.4042 0.3895 0.3767 0.3644 0.3564 $R^{2}$ 0.9637 0.9606 0.9705 0.9740 0.9755 0.9763 0.9764 0.9755 Sigmoid RMSE 1.5850 0.9691 1.5702 1.2545 1.2594 1.1208 0.9594 1.9734 MAE 1.1870 0.7475 1.2129 0.9506 0.9707 0.8719 0.7301 1.6113 $R^{2}$ 0.7468 0.9181 0.8022 0.8799 0.8832 0.9097 0.9348 0.7250 Cosine RMSE 5.2944 5.8910 6.0766 6.1494 6.1255 6.0142 5.8459 5.6488 MAE 4.2685 4.8210 4.9456 4.9710 4.8940 4.7187 4.4806 4.2142 $R^{2}$1.8257 -2.0275 -1.9619 -1.8855 -1.7633 -1.5999 -1.4225 -1.2535 Training Linear RMSE 5.2967 5.9302 6.1264 6.2033 6.1787 6.0624 5.8818 5.6703 MAE 4.2821 4.8539 4.9820 5.0092 4.9391 4.7727 4.5383 4.2713 $R^{2}$1.7817 -2.0309 -1.9525 -1.8691 -1.7526 -1.6012 -1.4342 -1.2712 Polynomial RMSE 0.9698 0.7981 0.8325 0.8109 0.7800 0.8850 0.9109 0.8821 MAE 0.7195 0.6049 0.6223 0.6265 0.6016 0.6786 0.6957 0.6847 $R^{2}$ 0.9067 0.9451 0.9455 0.9510 0.9561 0.9446 0.9416 0.9450 RBF RMSE 0.0333 0.2315 0.1039 0.4314 0.3995 0.3123 0.2316 0.1623 MAE 0.0183 0.1548 0.0602 0.3035 0.2726 0.2113 0.1520 0.1017 $R^{2}$ 0.9999 0.9954 0.9992 0.9861 0.9885 0.9931 0.9962 0.9981 Laplacian RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 Sigmoid RMSE 1.6486 0.7634 1.6019 1.0965 1.1632 0.9449 0.7869 2.0626 MAE 1.2481 0.5958 1.2083 0.8516 0.8922 0.7384 0.6158 1.6549 $R^{2}$ 0.7305 0.9498 0.7981 0.9104 0.9024 0.9368 0.9564 0.6995 Cosine RMSE 5.2979 5.9310 6.1279 6.2054 6.1808 6.0647 5.8841 5.6725 MAE 4.2842 4.8553 4.9839 5.0095 4.9333 4.7599 4.5186 4.2461 $R^{2}$1.7830 -2.0316 -1.9539 -1.8711 -1.7545 -1.6031 -1.4362 -1.2730
@@
Table XI: Learning performance metrics for the models trained with the hyperparameters listed in Table IV for $\mathscr{O}_{Sum}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{3}$ with labels determined by $\mathscr{O}_{Sum}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 11, though not all of them are visible within the chosen plot range.
C.7 Results for $\mathscr{O}_{XZ}$ and $\mathscr{H}_{4}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 0.8851 0.9240 0.9337 0.9401 0.9424 0.9408 0.9360 0.9293 MAE 0.8375 0.8844 0.8976 0.9061 0.9084 0.9053 0.8982 0.8878 $R^{2}$5.4877 -9.7666 -12.2134 -13.2260 -13.3083 -12.7522 -11.8802 -10.6947 Polynomial RMSE 0.1892 0.1754 0.1597 0.1425 0.1385 0.1373 0.1195 0.1270 MAE 0.1060 0.0914 0.0822 0.0782 0.0745 0.0814 0.0690 0.0742 $R^{2}$ 0.7034 0.6119 0.6135 0.6729 0.6911 0.7070 0.7900 0.7817 RBF RMSE 0.1407 0.1499 0.1218 0.1033 0.1232 0.1083 0.0950 0.0888 MAE 0.0724 0.0731 0.0522 0.0498 0.0650 0.0577 0.0507 0.0468 $R^{2}$ 0.8360 0.7167 0.7751 0.8281 0.7555 0.8177 0.8674 0.8931 Laplacian RMSE 0.0974 0.1049 0.1282 0.1187 0.1159 0.1094 0.1003 0.0934 MAE 0.0514 0.0521 0.0715 0.0686 0.0658 0.0623 0.0575 0.0536 $R^{2}$ 0.9214 0.8612 0.7511 0.7732 0.7835 0.8140 0.8521 0.8817 Sigmoid RMSE 0.1908 0.1774 0.1569 0.1764 0.1635 0.5520 0.2105 29.3142 MAE 0.0981 0.0940 0.0815 0.1014 0.1055 0.3001 0.1254 17.0978 $R^{2}$ 0.6985 0.6030 0.6269 0.4990 0.5695 -3.7352 0.3487 -11636.6066 Cosine RMSE 0.8851 0.9218 0.9336 0.9391 0.9407 0.9384 0.9332 0.9261 MAE 0.8376 0.8821 0.8973 0.9050 0.9068 0.9033 0.8959 0.8851 $R^{2}$5.4872 -9.7161 -12.2097 -13.1944 -13.2552 -12.6843 -11.8028 -10.6139 Training Linear RMSE 0.8912 0.9224 0.9317 0.9352 0.9363 0.9345 0.9300 0.9233 MAE 0.8470 0.8860 0.8965 0.9001 0.9002 0.8968 0.8903 0.8810 $R^{2}$6.5416 -11.1825 -12.0591 -12.1214 -11.7599 -11.1197 -10.3358 -9.4707 Polynomial RMSE 0.1576 0.1181 0.1125 0.1046 0.1113 0.1131 0.0909 0.0955 MAE 0.0884 0.0656 0.0616 0.0616 0.0641 0.0704 0.0514 0.0555 $R^{2}$ 0.7641 0.8003 0.8095 0.8358 0.8196 0.8226 0.8916 0.8879 RBF RMSE 0.0812 0.0831 0.0444 0.0221 0.0914 0.0796 0.0652 0.0521 MAE 0.0434 0.0440 0.0226 0.0103 0.0510 0.0429 0.0349 0.0275 $R^{2}$ 0.9374 0.9012 0.9703 0.9927 0.8783 0.9120 0.9443 0.9666 Laplacian RMSE 0.0000 0.0000 0.0076 0.0048 0.0034 0.0026 0.0019 0.0002 MAE 0.0000 0.0000 0.0051 0.0032 0.0022 0.0016 0.0012 0.0001 $R^{2}$ 1.0000 1.0000 0.9991 0.9996 0.9998 0.9999 1.0000 1.0000 Sigmoid RMSE 0.1130 0.1185 0.0985 0.1668 0.1557 0.6193 0.2038 32.9478 MAE 0.0658 0.0673 0.0550 0.0976 0.0989 0.3556 0.1172 20.1928 $R^{2}$ 0.8788 0.7988 0.8540 0.5826 0.6472 -4.3227 0.4554 -13333.5415 Cosine RMSE 0.8922 0.9237 0.9325 0.9363 0.9374 0.9357 0.9311 0.9243 MAE 0.8483 0.8874 0.8969 0.9003 0.9003 0.8967 0.8896 0.8797 $R^{2}$6.5589 -11.2173 -12.0804 -12.1504 -$1$11.1499 -10.3634 -9.4954
@@
Table XII: Learning performance metrics for the models trained with the hyperparameters listed in Table V for $\mathscr{O}_{XZ}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{4}$ with labels determined by $\mathscr{O}_{XZ}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 13, though not all of them are visible within the chosen plot range.
C.8 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{4}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear RMSE 7.8717 8.2751 8.3789 8.4404 8.4590 8.4367 8.3851 8.3171 MAE 7.3590 7.8947 8.0577 8.1362 8.1498 8.1097 8.0297 7.9231 $R^{2}$6.0930 -10.4791 -12.5001 -13.3118 -13.1500 -12.4112 -11.3274 -10.0421 Polynomial RMSE 1.1086 1.3555 1.2815 1.2502 1.4368 1.1573 1.1370 1.1639 MAE 0.7069 0.7836 0.6983 0.7379 0.8742 0.6691 0.6764 0.7258 $R^{2}$ 0.8593 0.6920 0.6842 0.6860 0.5918 0.7476 0.7734 0.7838 RBF RMSE 0.8913 1.0930 0.8644 0.7940 1.0286 0.8982 0.8055 0.7614 MAE 0.5528 0.5934 0.4079 0.3943 0.5672 0.5084 0.4461 0.4183 $R^{2}$ 0.9091 0.7997 0.8563 0.8734 0.7908 0.8480 0.8862 0.9075 Laplacian RMSE 0.5859 0.7574 0.9806 0.9716 0.9516 0.8828 0.8107 0.7550 MAE 0.3555 0.4011 0.5750 0.5742 0.5497 0.5137 0.4757 0.4492 $R^{2}$ 0.9607 0.9038 0.8151 0.8103 0.8209 0.8532 0.8848 0.9090 Sigmoid RMSE 1.1541 1.3469 1.1936 1.5142 1.4086 5.2078 1.9184 223.0957 MAE 0.7069 0.7720 0.6515 0.8967 0.9337 2.8464 1.1720 130.1390 $R^{2}$ 0.8475 0.6959 0.7261 0.5394 0.6076 -4.1101 0.3548 -7943.9105 Cosine RMSE 7.8679 8.2535 8.3761 8.4298 8.4419 8.4145 8.3585 8.2868 MAE 7.3541 7.8779 8.0550 8.1268 8.1352 8.0915 8.0086 7.8983 $R^{2}$6.0862 -10.4191 -12.4909 -13.2760 -13.0931 -12.3407 -11.2493 -9.9617 Training Linear RMSE 7.9094 8.2592 8.3579 8.3957 8.4016 8.3794 8.3295 8.2601 MAE 7.4386 7.9118 8.0255 8.0652 8.0626 8.0271 7.9567 7.8595 $R^{2}$6.6620 -10.8265 -11.6110 -11.5868 -11.1922 -10.5742 -9.7730 -8.8826 Polynomial RMSE 0.7762 0.9656 0.9374 0.9621 1.1591 0.9408 0.9154 0.9165 MAE 0.4871 0.5608 0.5223 0.5788 0.6837 0.5476 0.5393 0.5459 $R^{2}$ 0.9262 0.8383 0.8414 0.8347 0.7679 0.8541 0.8699 0.8783 RBF RMSE 0.5106 0.6467 0.3556 0.1823 0.7904 0.6843 0.5594 0.4568 MAE 0.3166 0.3636 0.1849 0.0864 0.4471 0.3794 0.3089 0.2443 $R^{2}$ 0.9681 0.9275 0.9772 0.9941 0.8921 0.9228 0.9514 0.9698 Laplacian RMSE 0.0000 0.0000 0.0602 0.0398 0.0281 0.0204 0.0000 0.0000 MAE 0.0000 0.0000 0.0412 0.0266 0.0184 0.0132 0.0000 0.0000 $R^{2}$ 1.0000 1.0000 0.9993 0.9997 0.9999 0.9999 1.0000 1.0000 Sigmoid RMSE 0.7593 0.9347 0.8167 1.4820 1.3557 5.8739 1.8615 250.6840 MAE 0.4922 0.5471 0.4609 0.8870 0.8745 3.3699 1.0980 153.6522 $R^{2}$ 0.9294 0.8485 0.8796 0.6078 0.6826 -4.6875 0.4620 -9101.2820 Cosine RMSE 7.9167 8.2705 8.3644 8.4045 8.4116 8.3895 8.3394 8.2697 MAE 7.4446 7.9173 8.0274 8.0670 8.0634 8.0236 7.9482 7.8462 $R^{2}$6.6762 -10.8590 -11.6306 -11.6133 -$1$10.6022 -9.7986 -8.9055
@@
Table XIII: Learning performance metrics for the models trained with the hyperparameters listed in Table V for $\mathscr{O}_{Sum}$. The metrics are calculated based on the predictions made by the models on the training and testing sets for $\mathscr{H}_{4}$ with labels determined by $\mathscr{O}_{Sum}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets. All values in this table are plotted in Figure 13, though not all of them are visible within the chosen plot range.

### v2

@@ D.5 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{1}$
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$2.6121 -3.7405 -3.5677 -3.5157 -3.3017 -3.1259 -2.8970 -2.6603 RMSE 6.6797 7.2117 7.2580 7.3482 7.2775 7.2564 7.1498 7.0573 MAE 5.8351 6.4148 6.4226 6.5026 6.3839 6.3326 6.1876 6.0515 Polynomial $R^{2}$ 0.9305 0.8697 0.8734 0.8188 0.8258 0.8168 0.8246 0.8211 RMSE 0.9266 1.1954 1.2084 1.4719 1.4644 1.5292 1.5170 1.5602 MAE 0.6546 0.8219 0.8558 1.0416 1.0461 1.0767 1.0694 1.1116 RBF $R^{2}$ 0.9556 0.9289 0.9315 0.8979 0.8985 0.9094 0.9208 0.9283 RMSE 0.7402 0.8833 0.8885 1.1049 1.1179 1.0755 1.0196 0.9878 MAE 0.5409 0.4966 0.4385 0.6916 0.6624 0.6367 0.5657 0.5190 Laplacian $R^{2}$ 0.9621 0.9652 0.9578 0.9555 0.9532 0.9429 0.9503 0.9596 RMSE 0.6841 0.6176 0.6979 0.7291 0.7592 0.8534 0.8078 0.7417 MAE 0.4643 0.3579 0.3771 0.3990 0.4168 0.5653 0.5250 0.4746 Sigmoid $R^{2}$ 0.7472 0.8715 0.6827 0.8175 0.8290 0.8526 0.6879 0.6413 RMSE 1.7673 1.1875 1.9128 1.4774 1.4508 1.3716 2.0234 2.2092 MAE 1.3062 0.7939 1.4378 1.0354 1.0400 0.9645 1.5305 1.7456 Cosine $R^{2}$2.6387 -3.7354 -3.5476 -3.4864 -3.2995 -3.1673 -2.9321 -2.6887 RMSE 6.7042 7.2079 7.2421 7.3244 7.2756 7.2927 7.1820 7.0846 MAE 5.8541 6.4031 6.3958 6.4589 6.3751 6.3714 6.2191 6.0739 Training Linear $R^{2}$2.1828 -3.1641 -3.0659 -2.9840 -2.8636 -2.6929 -2.5129 -2.3154 RMSE 6.4890 7.0385 7.0937 7.1710 7.1516 7.1017 6.9922 6.8715 MAE 5.5720 6.1532 6.1925 6.2540 6.1792 6.1015 5.9658 5.8112 Polynomial $R^{2}$ 0.9596 0.9364 0.9343 0.8788 0.8862 0.8801 0.8902 0.8845 RMSE 0.7307 0.8701 0.9015 1.2510 1.2271 1.2795 1.2360 1.2827 MAE 0.5112 0.6282 0.6547 0.8919 0.8785 0.8960 0.8745 0.9044 RBF $R^{2}$ 0.9753 0.9822 0.9990 0.9603 0.9754 0.9715 0.9825 0.9880 RMSE 0.5713 0.4603 0.1088 0.7159 0.5706 0.6236 0.4932 0.4137 MAE 0.4250 0.3093 0.0643 0.4913 0.3904 0.4216 0.3307 0.2663 Laplacian $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 0.9997 0.9998 0.9999 RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0648 0.0492 0.0397 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0463 0.0352 0.0279 Sigmoid $R^{2}$ 0.7473 0.9429 0.7354 0.8782 0.8768 0.9162 0.7237 0.6295 RMSE 1.8283 0.8245 1.8096 1.2541 1.2770 1.0701 1.9611 2.2970 MAE 1.3103 0.5973 1.3112 0.9025 0.9082 0.7856 1.4166 1.7306 Cosine $R^{2}$2.1793 -3.1696 -3.0751 -2.9934 -2.8659 -2.6827 -2.5044 -2.3079 RMSE 6.4854 7.0431 7.1016 7.1794 7.1538 7.0918 6.9837 6.8637 MAE 5.5633 6.1464 6.1758 6.2249 6.1695 6.1030 5.9616 5.8011
@@
Table D.5: ($\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.2. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.6 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{2}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$2.4632 -4.3090 -5.1549 -5.2246 -$5$4.7658 -4.5452 -4.3089 RMSE 6.6084 6.9384 7.1717 7.2117 7.2634 7.2450 7.2467 7.2143 MAE 5.7357 6.2669 6.5641 6.6096 6.6403 6.5921 6.5668 6.5059 Polynomial $R^{2}$ 0.8230 0.8447 0.7799 0.8029 0.7957 0.8045 0.6727 0.7990 RMSE 1.4942 1.1867 1.3563 1.2833 1.3342 1.3341 1.7604 1.4036 MAE 1.1004 0.8443 1.0029 0.9740 1.0096 1.0189 1.2957 1.0740 RBF $R^{2}$ 0.9706 0.9049 0.9022 0.8957 0.8967 0.8965 0.9053 0.9156 RMSE 0.6089 0.9284 0.9039 0.9337 0.9487 0.9708 0.9471 0.9099 MAE 0.3170 0.6004 0.5720 0.6676 0.6700 0.6674 0.6390 0.6063 Laplacian $R^{2}$ 0.9098 0.8973 0.8990 0.9088 0.9108 0.8862 0.8865 0.8897 RMSE 1.0664 0.9648 0.9189 0.8730 0.8813 1.0177 1.0370 1.0398 MAE 0.7712 0.6657 0.6309 0.6007 0.5915 0.7304 0.7410 0.7411 Sigmoid $R^{2}$ 0.4967 0.8416 0.5619 0.8142 0.8178 0.8291 0.6329 0.5820 RMSE 2.5193 1.1984 1.9133 1.2459 1.2601 1.2474 1.8645 2.0244 MAE 1.9383 0.8575 1.3799 0.9470 0.9429 0.8848 1.3602 1.4461 Cosine $R^{2}$2.4709 -4.3156 -5.1549 -5.2228 -5.0525 -4.7597 -4.5384 -4.2995 RMSE 6.6158 6.9427 7.1717 7.2107 7.2616 7.2411 7.2422 7.2080 MAE 5.7417 6.2682 6.5637 6.6062 6.6348 6.5832 6.5569 6.4933 Training Linear $R^{2}$2.5404 -4.3985 -5.3317 -5.2921 -5.1667 -4.8959 -4.6703 -4.4283 RMSE 6.6747 6.9451 7.2232 7.2249 7.2857 7.2512 7.2503 7.2167 MAE 5.7924 6.2699 6.6278 6.6248 6.6670 6.6053 6.5781 6.5165 Polynomial $R^{2}$ 0.8817 0.8715 0.7817 0.8027 0.7974 0.8080 0.7078 0.8052 RMSE 1.2201 1.0716 1.3412 1.2795 1.3206 1.3084 1.6460 1.3672 MAE 0.9328 0.7624 0.9493 0.9258 0.9539 0.9551 1.2068 1.0022 RBF $R^{2}$ 0.9999 0.9441 0.9964 0.9292 0.9275 0.9478 0.9626 0.9752 RMSE 0.0252 0.7066 0.1727 0.7663 0.7902 0.6823 0.5885 0.4879 MAE 0.0142 0.4788 0.1020 0.5265 0.5465 0.4763 0.4020 0.3227 Laplacian $R^{2}$ 0.9539 0.9736 0.9995 0.9998 1.0000 0.9860 0.9891 0.9912 RMSE 0.7618 0.4858 0.0639 0.0435 0.0000 0.3533 0.3186 0.2906 MAE 0.5515 0.3406 0.0450 0.0301 0.0000 0.2641 0.2397 0.2191 Sigmoid $R^{2}$ 0.5367 0.8811 0.5941 0.8180 0.8082 0.8815 0.6817 0.5779 RMSE 2.4146 1.0306 1.8288 1.2287 1.2850 1.0282 1.7178 2.0124 MAE 1.8768 0.7487 1.3039 0.8932 0.9275 0.7365 1.2508 1.4609 Cosine $R^{2}$2.5378 -4.4031 -5.3380 -5.3013 -5.1766 -4.9077 -4.6811 -4.4397 RMSE 6.6723 6.9480 7.2267 7.2301 7.2915 7.2585 7.2572 7.2243 MAE 5.7902 6.2753 6.6324 6.6318 6.6754 6.6159 6.5880 6.5268
@@
Table D.6: ($\mathscr{O}_{Sum}$ with $\mathscr{H}_{2}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.3. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{Sum}$ with $\mathscr{H}_{2}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.7 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{3}$
@@
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$1.8253 -2.0274 -1.9619 -1.8855 -1.7631 -1.5998 -1.4225 -1.2536 RMSE 5.2940 5.8909 6.0765 6.1494 6.1253 6.0140 5.8459 5.6489 MAE 4.2697 4.8224 4.9492 4.9773 4.9052 4.7378 4.5049 4.2420 Polynomial $R^{2}$ 0.8647 0.9194 0.9221 0.9244 0.9296 0.9222 0.9160 0.9140 RMSE 1.1585 0.9611 0.9857 0.9956 0.9778 1.0404 1.0888 1.1037 MAE 0.8594 0.7375 0.7432 0.7518 0.7466 0.7980 0.8239 0.8420 RBF $R^{2}$ 0.9852 0.9867 0.9882 0.9703 0.9776 0.9810 0.9836 0.9850 RMSE 0.3834 0.3908 0.3831 0.6243 0.5514 0.5142 0.4808 0.4614 MAE 0.2247 0.2704 0.2342 0.4376 0.3829 0.3485 0.3089 0.2831 Laplacian $R^{2}$ 0.9637 0.9606 0.9705 0.9740 0.9755 0.9763 0.9764 0.9755 RMSE 0.6000 0.6719 0.6060 0.5842 0.5770 0.5740 0.5768 0.5884 MAE 0.3753 0.4879 0.4285 0.4042 0.3895 0.3767 0.3644 0.3564 Sigmoid $R^{2}$ 0.7468 0.9181 0.8022 0.8799 0.8832 0.9097 0.9348 0.7250 RMSE 1.5850 0.9691 1.5702 1.2545 1.2594 1.1208 0.9594 1.9734 MAE 1.1870 0.7475 1.2129 0.9506 0.9707 0.8719 0.7301 1.6113 Cosine $R^{2}$1.8257 -2.0275 -1.9619 -1.8855 -1.7633 -1.5999 -1.4225 -1.2535 RMSE 5.2944 5.8910 6.0766 6.1494 6.1255 6.0142 5.8459 5.6488 MAE 4.2685 4.8210 4.9456 4.9710 4.8940 4.7187 4.4806 4.2142 Training Linear $R^{2}$1.7817 -2.0309 -1.9525 -1.8691 -1.7526 -1.6012 -1.4342 -1.2712 RMSE 5.2967 5.9302 6.1264 6.2033 6.1787 6.0624 5.8818 5.6703 MAE 4.2821 4.8539 4.9820 5.0092 4.9391 4.7727 4.5383 4.2713 Polynomial $R^{2}$ 0.9067 0.9451 0.9455 0.9510 0.9561 0.9446 0.9416 0.9450 RMSE 0.9698 0.7981 0.8325 0.8109 0.7800 0.8850 0.9109 0.8821 MAE 0.7195 0.6049 0.6223 0.6265 0.6016 0.6786 0.6957 0.6847 RBF $R^{2}$ 0.9999 0.9954 0.9992 0.9861 0.9885 0.9931 0.9962 0.9981 RMSE 0.0333 0.2315 0.1039 0.4314 0.3995 0.3123 0.2316 0.1623 MAE 0.0183 0.1548 0.0602 0.3035 0.2726 0.2113 0.1520 0.1017 Laplacian $R^{2}$ 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 RMSE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 MAE 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 Sigmoid $R^{2}$ 0.7305 0.9498 0.7981 0.9104 0.9024 0.9368 0.9564 0.6995 RMSE 1.6486 0.7634 1.6019 1.0965 1.1632 0.9449 0.7869 2.0626 MAE 1.2481 0.5958 1.2083 0.8516 0.8922 0.7384 0.6158 1.6549 Cosine $R^{2}$1.7830 -2.0316 -1.9539 -1.8711 -1.7545 -1.6031 -1.4362 -1.2730 RMSE 5.2979 5.9310 6.1279 6.2054 6.1808 6.0647 5.8841 5.6725 MAE 4.2842 4.8553 4.9839 5.0095 4.9333 4.7599 4.5186 4.2461
@@
Table D.7: ($\mathscr{O}_{Sum}$ with $\mathscr{H}_{3}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.4. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{Sum}$ with $\mathscr{H}_{3}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.
D.8 Results for $\mathscr{O}_{Sum}$ and $\mathscr{H}_{4}$
Number of qubits $(n)$ Dataset Kernel Metric 5 10 15 20 25 30 35 40 Testing Linear $R^{2}$6.0930 -10.4791 -12.5001 -13.3118 -13.1500 -12.4112 -11.3274 -10.0421 RMSE 7.8717 8.2751 8.3789 8.4404 8.4590 8.4367 8.3851 8.3171 MAE 7.3590 7.8947 8.0577 8.1362 8.1498 8.1097 8.0297 7.9231 Polynomial $R^{2}$ 0.8593 0.6920 0.6842 0.6860 0.5918 0.7476 0.7734 0.7838 RMSE 1.1086 1.3555 1.2815 1.2502 1.4368 1.1573 1.1370 1.1639 MAE 0.7069 0.7836 0.6983 0.7379 0.8742 0.6691 0.6764 0.7258 RBF $R^{2}$ 0.9091 0.7997 0.8563 0.8734 0.7908 0.8480 0.8862 0.9075 RMSE 0.8913 1.0930 0.8644 0.7940 1.0286 0.8982 0.8055 0.7614 MAE 0.5528 0.5934 0.4079 0.3943 0.5672 0.5084 0.4461 0.4183 Laplacian $R^{2}$ 0.9607 0.9038 0.8151 0.8103 0.8209 0.8532 0.8848 0.9090 RMSE 0.5859 0.7574 0.9806 0.9716 0.9516 0.8828 0.8107 0.7550 MAE 0.3555 0.4011 0.5750 0.5742 0.5497 0.5137 0.4757 0.4492 Sigmoid $R^{2}$ 0.8475 0.6959 0.7261 0.5394 0.6076 -4.1101 0.3548 -7943.9105 RMSE 1.1541 1.3469 1.1936 1.5142 1.4086 5.2078 1.9184 223.0957 MAE 0.7069 0.7720 0.6515 0.8967 0.9337 2.8464 1.1720 130.1390 Cosine $R^{2}$6.0862 -10.4191 -12.4909 -13.2760 -13.0931 -12.3407 -11.2493 -9.9617 RMSE 7.8679 8.2535 8.3761 8.4298 8.4419 8.4145 8.3585 8.2868 MAE 7.3541 7.8779 8.0550 8.1268 8.1352 8.0915 8.0086 7.8983 Training Linear $R^{2}$6.6620 -10.8265 -11.6110 -11.5868 -11.1922 -10.5742 -9.7730 -8.8826 RMSE 7.9094 8.2592 8.3579 8.3957 8.4016 8.3794 8.3295 8.2601 MAE 7.4386 7.9118 8.0255 8.0652 8.0626 8.0271 7.9567 7.8595 Polynomial $R^{2}$ 0.9262 0.8383 0.8414 0.8347 0.7679 0.8541 0.8699 0.8783 RMSE 0.7762 0.9656 0.9374 0.9621 1.1591 0.9408 0.9154 0.9165 MAE 0.4871 0.5608 0.5223 0.5788 0.6837 0.5476 0.5393 0.5459 RBF $R^{2}$ 0.9681 0.9275 0.9772 0.9941 0.8921 0.9228 0.9514 0.9698 RMSE 0.5106 0.6467 0.3556 0.1823 0.7904 0.6843 0.5594 0.4568 MAE 0.3166 0.3636 0.1849 0.0864 0.4471 0.3794 0.3089 0.2443 Laplacian $R^{2}$ 1.0000 1.0000 0.9993 0.9997 0.9999 0.9999 1.0000 1.0000 RMSE 0.0000 0.0000 0.0602 0.0398 0.0281 0.0204 0.0000 0.0000 MAE 0.0000 0.0000 0.0412 0.0266 0.0184 0.0132 0.0000 0.0000 Sigmoid $R^{2}$ 0.9294 0.8485 0.8796 0.6078 0.6826 -4.6875 0.4620 -9101.2820 RMSE 0.7593 0.9347 0.8167 1.4820 1.3557 5.8739 1.8615 250.6840 MAE 0.4922 0.5471 0.4609 0.8870 0.8745 3.3699 1.0980 153.6522 Cosine $R^{2}$6.6762 -10.8590 -11.6306 -11.6133 -$1$10.6022 -9.7986 -8.9055 RMSE 7.9167 8.2705 8.3644 8.4045 8.4116 8.3895 8.3394 8.2697 MAE 7.4446 7.9173 8.0274 8.0670 8.0634 8.0236 7.9482 7.8462
Table D.8: ($\mathscr{O}_{Sum}$ with $\mathscr{H}_{4}$) Learning performance metrics for the models trained on all 1000 training datapoints using the hyperparameters listed in Table C.5. The metrics are calculated based on the predictions made by the models on the training and testing sets with labels determined by $\mathscr{O}_{Sum}$ with $\mathscr{H}_{4}$. The best values for each qubit number and each metric are in bold text for the training and testing datasets.